Water is flowing through a circular tube with a cross-sectional area 0.4m^2 at a rate of 8m/s.

Step 1: the cross-section of the tube gradually decreases to 0.2m^2,

Step 2: then the water shoots out upright from the tape (in the up vertical direction), and gains 4m in height.

What is the new velocity of the water after it gains 4m in height? [save to 1 decimal places]

Assume g=10m/s2; for step 1 and 2, the water is under the same atmosphere pressure;

Tips: Step 1 where the conservation of mass is applicable;

Step 2 where the conservation of energy is applicable.
Answer #1

v1=8m/s

A1=0.4m2, A2=0.2m2

Use continuity equation,

A1v1= A2v2

v2=(A1/A2)v1=(0.4/0.2)*8=16m/s

Now use Bernoulli’s principle

Pi+1/2 *ρvi2 +ρghi = Pf+1/2 *ρvf2 +ρghf

Pi= Pf

1/2 *ρvi2 +ρghi = Pf+1/2 *ρvf2 +ρghf

½ρ(vi2 – vf2) = ρg(hf - hi)

(vi2 – vf2) = 2g(hf - hi)

vf2= vi2 –2g(hf - hi)

vf=sqrt[vi2 –2g(hf - hi)]

Plug values, vi= 16m/s, hf - hi=4m

vf=sqrt[162 –2*10*10]

vf=7.5 m/s